OpenStudy (anonymous):

An elementary student kicks a ball straight into the air with a velocity of 16 feet/sec. If acceleration due to gravity is −32 ft/sec2, how many seconds after it leaves his foot will it take the ball to reach its highest point? Assume the position at time t = 0 is 0 feet. a: 1.25 b: 6 c: 0.50 d: Cannot be determined

OpenStudy (anonymous):

what is the formula you have to use?

OpenStudy (anonymous):

unprovided :(

OpenStudy (primeralph):

v = u+gt at highest point, v = 0 so 0 = 16-32t Solve for t

OpenStudy (anonymous):

2 :) thank u!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

not an option

OpenStudy (primeralph):

It's not 2.

OpenStudy (anonymous):

-16=-32t divide by -32 on both before i erase this i caught my mistake. 0.5!!!!!!! thanks :)

OpenStudy (primeralph):

Yeah.